# Nonlocal Contributions to the Turbulent Cascade in Magnetohydrodynamic Plasmas

### 19 Mar 2024 v2

@@ I Introduction
The physical mechanisms underlying the solar wind - a continuous flow of charged particles emitted from the solar corona - have been hypothesized since the beginning of the past century. Based on Biermann’s observations of comet tail motion relative to the Sun [1], Parker provided the first comprehensive theoretical description of the solar wind as a rapidly expanding outer coronal atmosphere overcoming the Sun’s gravitational field [2]. Assuming an adiabatic solar wind expansion, Parker also considered the radial temperature profile $T(r)$ as a function of the heliocentric distance $r$. Nonetheless, a purely adiabatic expansion of the solar wind, which suggests a radial temperature profile of proton temperature as $T(r)\sim r^{-4/3}$, differs from actual solar wind plasma measurements that exhibit a much slower decay with increasing distances $r$ [3, 4, 5].
A possible mechanism that would supply heat to the expanding solar wind - and thus explain the slower decay of the temperature profile - is the dissipation of turbulent kinetic and magnetic energy into heat predicted by a phenomenological description of magnetohydrodynamic (MHD) turbulence [6] (we also refer the reader to the recent review article on scaling laws in solar wind turbulence [7]). The concept of turbulent cascades, which has been put forth in the context of hydrodynamic turbulence by Kolmogorov [8], Heisenberg [9], von Weizsäcker [10], and Onsager [11], highlights the fact that turbulent motions are essentially transport processes of energy in scale. The existence of energy cascades in fluid and plasma turbulence has been assessed experimentally (e.g., in channel flows and fusion plasmas [12, 13]), as well as by atmospheric [14], oceanic [15], and space plasma measurements [16]. A central quantity in phenomenological descriptions of such cascade processes is the averaged energy dissipation rate $\langle\varepsilon_{kin}\rangle$, which is assumed to entirely characterize the transfer of energy from the injection scale, where the turbulent flow is stirred, to small scales where energy is dissipated into heat.
@@
where ${\bf u}({\bf x},t)$ denotes the velocity field, ${\bf h}({\bf x},t)=\sqrt{\mu/4\pi\rho}\;{\bf H}({\bf x},t)$ the rescaled magnetic field, $\rho$ the density of the fluid, $\mu$ the permeability, $\nu$ the kinematic viscosity, and $\lambda$ the magnetic diffusivity (see Appendix A for further details). Fig. 1 (a) depicts a two-dimensional cut through the total energy dissipation rate field $\varepsilon_{tot}({\bf x},t)=\varepsilon_{kin}({\bf x},t)+\varepsilon_{mag}({\bf x},t)$ from a direct numerical simulation (DNS) of MHD turbulence (see Tab. 1 for further details). The dissipation field is highly fluctuating and exhibits strong gradients organized in current sheets in contrast to hydrodynamic turbulence [17]; see Fig. 1 (b), where the peaks are more localized, hinting at the presence of vortex tubes [18, 19].
The solar wind is weakly collisional, which implies that energy cascades from the MHD scales further down to proton and electron scales [20]. Due to the limited resolution of the instruments onboard most spacecraft and the presence of measurement noise, strong gradients of the energy dissipation at kinetic scales can typically not be resolved directly from in situ solar wind observations. Nonetheless, the scale-wise energy transfer at the scales at which the interplanetary plasma has a fluid-like behavior can be estimated indirectly from third-order moments of velocity and magnetic fields [21, 7].
@@
This article presents new evidence for nonlocal contributions to kinetic and magnetic energy transfer. To this end, we first revisit the derivation of the P&P law [23] in Sec. II, obtaining from first principles an additional nonlocal term that accounts for the influence of magnetic fields induced at large scales on fluctuations at small scales as suggested by the Iroshnikov-Kraichnan phenomenology. The importance of this nonlocal term for estimations of turbulent energy transfer based on third-order moments is then assessed by DNS of homogeneous and isotropic MHD turbulence in Sec. III, and the validity of the model proposed is demonstrated in the plasma under study. Although the original P&P law has been verified in numerous DNS of 2D and 3D MHD turbulence [54, 55, 56] (see also [7] for further references), here, we provide analytical and numerical evidence that there can be plasma configurations for which considering the nonlocal contributions to the MHD cascade is critical to correctly infer the plasma heating rate through the third-order law approach. These configurations can be representative of heliospheric plasmas, whose highly dynamical state is such that solar wind may at times be characterized by enhanced turbulence and nonlocal couplings of kinetic and magnetic modes, due to magnetic reconnection [57], magnetic switchbacks [58, 59] or the interaction between fast and slow solar wind streams in the ecliptic [40].
@@ II Revisiting the P&P law using the invariant theory of MHD turbulence
where we imply summation over identical indices and introduce the root mean square velocity and magnetic fields (please see Appendix A for further details on the derivation of Eq. (4)). In his seminal work, Chandrasekhar generalized the work of von Kármán and Howarth [61] to an invariant theory of homogeneous and isotropic MHD turbulence [60]. Here, we consider evolution equations for the velocity and magnetic field correlation tensors, $C_{i\;j}^{\mathbf{u}\mathbf{u}}({\bf r},t)=\left\langle u_{i}u_{j}^{\prime}\right\rangle=\left\langle u_{i}({\bf x},t)u_{j}({\bf x}+{\bf r},t)\right\rangle$ and $C_{i\;j}^{\mathbf{h}\mathbf{h}}({\bf r},t)=\left\langle h_{i}h_{j}^{\prime}\right\rangle=\left\langle h_{i}({\bf x},t)h_{j}({\bf x}+{\bf r},t)\right\rangle$, which can be derived from the MHD equations (under the assumption of homogeneity, see Appendix B and C) according to
@@
$A_{[ik]j}^{\mathbf{u}\mathbf{h}{\mathbf{h}}}({\bf r},t)=$ $\left\langle(h_{i}u_{k}-h_{k}u_{i})h_{j}^{\prime}\right\rangle$
@@
Furthermore, the indices $r$ and $t$ denote the longitudinal and transverse projections, ${\bf u}_{r}=\frac{{\bf r}}{r}\left(\frac{{\bf r}}{r}\cdot{\bf u}\right)$ and ${\bf u}_{t}=-\left(\frac{{\bf r}}{r}\times\left(\frac{{\bf r}}{r}\times{\bf u}\right)\right)$ respectively (the same projections hold for the magnetic field ${\bf h}({\bf x},t)$, see Appendix F for further derivation). Following Chandrasekhar [60] (see Appendix C for further derivations), we obtain evolution equations for the longitudinal velocity and magnetic field correlation functions
@@
$\frac{\partial}{\partial t}C_{r\;r}^{\mathbf{h}\mathbf{h}}(r,t)=$ $\frac{4}{r}A_{[rt]t}^{\mathbf{u}\mathbf{h}\mathbf{h}}(r,t)+2\lambda\frac{1}{r^{4}}\frac{\partial}{\partial r}r^{4}\frac{\partial}{\partial r}C_{r\;r}^{\mathbf{h}\mathbf{h}}(r,t)\;.$ (9)
Eq. (8) is the generalization of the von Kármán-Howarth equation to MHD turbulence and the additional third-order correlation is due to the Lorentz force $C_{r\;r\;r}^{\mathbf{h}\mathbf{h}\mathbf{u}}(r,t)$ in the MHD equations. A similar equation also holds for the second-order longitudinal magnetic field correlation function (9) where the third-order correlation is due to the advection by the velocity field and the corresponding transformation of the electric field ${\bf E}^{\prime}={\bf E}+\mu{\bf u}\times{\bf H}$. Similar equations were derived recently in the context of Hall MHD [62], where the Hall term modifies such a transformation.
@@
$\left.\left.+\frac{\partial}{\partial r}\left[\nu S_{r\;r}^{\mathbf{u}\mathbf{u}}(r,t)+\lambda S_{r\;r}^{\mathbf{h}\mathbf{h}}(r,t)\right]\right)\right]-\frac{4}{r}A_{[rt]t}^{\mathbf{u}\mathbf{h}\mathbf{h}}(r,t)\;.$ (17)
@@
In the inertial range, i.e., for $r$ smaller than integral- and larger than dissipation-length scales, we neglect the viscous terms in square brackets and obtain a relation between third-order statistics on the l.h.s. and a term that is proportional to $r$ on the r.h.s. whose magnitude is determined by the total averaged local energy dissipation rate $\langle\varepsilon_{tot}\rangle$ [63]. Nonetheless, in contrast to the hydrodynamic case (3) and to the original derivation by Politano and Pouquet [23, 24] where the energy transfer is purely local in scale $r$, we obtain a scaling law in the MHD inertial range (18) that exhibits a nonlocal term stemming from the magnetic induction equation and is a direct consequence of the anti-symmetric tensorial form (7). As shown in Appendix E, we can re-cast this nonlocal term as
@@
which thus represents the influence of magnetic field structures induced at larger scales $r^{\prime}$ on the local fluctuations at scale $r$ in the sense of the Alfvén effect [27, 26, 28, 29]. We will now consider further implications of this additional nonlocality for the conservation of energy at small scales and the phenomenon of solar wind heating by turbulent cascades. In more detail, for small scales $r$ in the vicinity of the dissipation range, nonlinear transfer terms on the l.h.s. in Eq. (18) should decay faster than the terms on the r.h.s. [22]. From a Taylor expansion around $r=0$, we obtain $S_{r\;r\,r}^{\mathbf{u}\mathbf{u}\mathbf{u}}(r)\sim r^{3}$ as well as
@@
Figure 3: (a) Assessment of the four-fifths law in MHD turbulence from DNS ($1024^{3}$) of forced MHD turbulence simulation. The dashed line represents the r.h.s. of Eq. (18, which is determined from the total energy dissipation rate $\varepsilon_{tot}$ of the snapshot. The top curve corresponds the original P&P law $S_{r\;r\;r}^{\mathbf{u}\mathbf{u}\mathbf{u}}(r)-6C_{r\;r\;r}^{\mathbf{h}\mathbf{h}\mathbf{u}}(r)$ derived in [23, 23] whereas the additional nonlocal term in Eq. (18) is closer to the dashed line. The mixed tensors verify the dissipation range behavior (22). The lowest curve represents the third-order longitudinal velocity structure function $S_{r\;r\;r}^{\mathbf{u}\mathbf{u}\mathbf{u}}(r)$. (b) Same as in (a) but compensated by the r.h.s. in Eq. (18). The original P&P law (top curve) overestimates the total energy dissipation rate $\varepsilon_{tot}$ by a factor of 1.71 and peaks in front of the inertial range, whereas the inclusion of the nonlocal term in Eq. (18) leads to the correct prediction (dashed line).
@@
$\left.\frac{\partial A_{[rt]r}^{\mathbf{u}\mathbf{h}\mathbf{h}}(r)}{\partial r}\right|_{r=0}=\frac{5}{4}\left.\frac{\partial C_{r\;r\;r}^{\mathbf{h}\mathbf{h}\mathbf{u}}(r)}{\partial r}\right|_{r=0}\;,$ (22)
@@ III Verification of the third-order law via direct numerical simulations of MHD turbulence
To verify Eq. (18), we performed DNS of homogeneous and isotropic MHD turbulence in a box of $1024^{3}$ grid points with characteristic turbulence parameters summarized in Tab. 1. The forcing scheme consists of a divergence-free random forcing [64] that is applied to the evolution equation of the velocity field (i.e., the evolution equation for the magnetic field is not actively forced). As shown in Fig. 2, after an initial phase, the forcing scheme leads to nearly constant kinetic and magnetic energy dissipation rates $\langle\varepsilon_{kin}\rangle$ and $\langle\varepsilon_{mag}\rangle$ over time. Fig. 3 depicts the third-order moments that enter Eq. (18). The lowest curve corresponds to the third-order velocity structure function $S_{r\;r\;r}^{\mathbf{u}\mathbf{u}\mathbf{u}}(r)$. Interestingly, this curve exhibits three distinct scaling regions: In the dissipation range, we observe the predicted scaling $\sim r^{3}$ followed by a region that scales in agreement with the Iroshnikov-Kraichnan phenomenology $\sim r^{3/4}$ whereas a linear increase seems to dominate at larger $r$. Furthermore, in the dissipation range, we can directly verify the relation between the symmetric and anti-symmetric correlations (22). In the case under study, the original prediction [23, 24] without the additional nonlocal source term overestimates the total energy dissipation rate $\langle\varepsilon_{tot}\rangle$ (dashed line) and does not vanish for small $r$, therefore, violating energy conservation. This becomes even more apparent in the compensated plot in Fig. 3 (b), which suggests that including the source term results in the correct scaling in the inertial range in between the largest Taylor scale $\lambda_{kin}$ and the smallest integral length scale $L_{mag}$. In contrast, the scaling obtained using the P&P law [23, 24] (green curve) would plateau before this inertial range and overestimate the $\langle\varepsilon_{tot}\rangle$ by a factor of 1.71. We must emphasize that this discrepancy between the original P&P law and (18) might have been barely detectable in previous numerical studies of MHD turbulence [54, 55, 56] due for instance to low magnetic and fluid Reynolds numbers and signatures of the large-scale anisotropic forcing scheme (e.g., by Taylor-Green vortices). Indeed, MHD and kinetic simulations reproduce plasma dynamics and features peculiar to the set of governing parameters defined for specific runs, which may differ significantly from actual solar wind parameters. Due to the insufficient computational power of today’s supercomputers, constraints in reproducing the solar wind numerically are technical but also inherent since solar wind dynamics develop in a vast parameter space, hardly reproducible even with large ensembles of simulations.

### 10 Apr 2024 v3

@@ I Introduction
The physical mechanisms underlying the solar wind - a continuous flow of charged particles emitted from the solar corona - have been hypothesized since the beginning of the past century. Based on Biermann’s observations of comet tail motion relative to the Sun Biermann (1951), Parker provided the first comprehensive theoretical description of the solar wind as a rapidly expanding outer coronal atmosphere overcoming the Sun’s gravitational field Parker (1958). Assuming an adiabatic solar wind expansion, Parker also considered the radial temperature profile $T(r)$ as a function of the heliocentric distance $r$. Nonetheless, a purely adiabatic expansion of the solar wind, which suggests a radial temperature profile of proton temperature as $T(r)\sim r^{-4/3}$, differs from actual solar wind plasma measurements that exhibit a much slower decay with increasing distances $r$ Matthaeus et al. (1999); Richardson and Smith (2003); Hundhausen (2012).
A possible mechanism that would supply heat to the expanding solar wind - and thus explain the slower decay of the temperature profile - is the dissipation of turbulent kinetic and magnetic energy into heat predicted by a phenomenological description of magnetohydrodynamic (MHD) turbulence Marino et al. (2008) (we also refer the reader to the recent review article on scaling laws in solar wind turbulence Marino and Sorriso-Valvo (2023)). The concept of turbulent cascades, which has been put forth in the context of hydrodynamic turbulence by Kolmogorov Kolmogorov (1941), Heisenberg Heisenberg (1948), von Weizsäcker v. Weizsäcker (1948), and Onsager Onsager (1949), highlights the fact that turbulent motions are essentially transport processes of energy in scale. The existence of energy cascades in fluid and plasma turbulence has been assessed experimentally (e.g., in channel flows and fusion plasmas Danaila et al. (2001); Conway (2008)), as well as by atmospheric Vierinen et al. (2019), oceanic Balwada et al. (2022), and space plasma measurements Telloni (2023). A central quantity in phenomenological descriptions of such cascade processes is the averaged energy dissipation rate $\langle\varepsilon_{kin}\rangle$, which is assumed to entirely characterize the transfer of energy from the injection scale, where the turbulent flow is stirred, to small scales where energy is dissipated into heat.
@@
where ${\bf u}({\bf x},t)$ denotes the velocity field, ${\bf h}({\bf x},t)=\sqrt{\mu/4\pi\rho}\;{\bf H}({\bf x},t)$ the rescaled magnetic field, $\rho$ the density of the fluid, $\mu$ the permeability, $\nu$ the kinematic viscosity, and $\lambda$ the magnetic diffusivity (see Appendix A for further details). Fig. 1 (a) depicts a two-dimensional cut through the total energy dissipation rate field $\varepsilon_{tot}({\bf x},t)=\varepsilon_{kin}({\bf x},t)+\varepsilon_{mag}({\bf x},t)$ from a direct numerical simulation (DNS) of MHD turbulence (see Tab. 1 for further details). The dissipation field is highly fluctuating and exhibits strong gradients organized in current sheets in contrast to hydrodynamic turbulence Marino et al. (2022); see Fig. 1 (b), where the peaks are more localized, hinting at the presence of vortex tubes Ishihara et al. (2009); Tanaka and Kida (1993).
The solar wind is weakly collisional, which implies that energy cascades from the MHD scales further down to proton and electron scales Matthaeus et al. (2020). Due to the limited resolution of the instruments onboard most spacecraft and the presence of measurement noise, strong gradients of the energy dissipation at kinetic scales can typically not be resolved directly from in situ solar wind observations. Nonetheless, the scale-wise energy transfer at the scales at which the interplanetary plasma has a fluid-like behavior can be estimated indirectly from third-order moments of velocity and magnetic fields Carbone et al. (2009); Marino and Sorriso-Valvo (2023).
@@
This article presents new evidence for nonlocal contributions to kinetic and magnetic energy transfer. To this end, we first revisit the derivation of the P&P law Politano and Pouquet (1998a) in Sec. II, obtaining from first principles an additional nonlocal term that accounts for the influence of magnetic fields induced at large scales on fluctuations at small scales as suggested by the Iroshnikov-Kraichnan phenomenology. The importance of this nonlocal term for estimations of turbulent energy transfer based on third-order moments is then assessed by DNS of homogeneous and isotropic MHD turbulence in Sec. III, and the validity of the model proposed is demonstrated in the plasma under study. Although the original P&P law has been verified in numerous DNS of two- and three-dimensional MHD turbulence Mininni and Pouquet (2009); Sorriso-Valvo et al. (2002); Yousef et al. (2007) (see also Marino and Sorriso-Valvo (2023) for further references), here, we provide analytical and numerical evidence that there can be plasma configurations for which considering the nonlocal contributions to the MHD cascade is critical to correctly infer the plasma heating rate through the third-order law approach. These configurations can be representative of heliospheric plasmas, whose highly dynamical state is such that solar wind may at times be characterized by enhanced turbulence and nonlocal couplings of kinetic and magnetic modes, due to magnetic reconnection Cerri et al. (2022), magnetic switchbacks Hernandez et al. (2021); Telloni et al. (2022) or the interaction between fast and slow solar wind streams in the ecliptic Marino et al. (2011).
@@ II Revisiting the P&P law using the invariant theory of MHD turbulence
where we imply summation over identical indices and introduce the root mean square velocity and magnetic fields (please see Appendix A for further details on the derivation of Eq. (4)). In his seminal work, Chandrasekhar generalized the work of von Kármán and Howarth de Karman and Howarth (1938) to an invariant theory of homogeneous and isotropic MHD turbulence Chandrasekhar (1951). Here, we consider evolution equations for the velocity and magnetic field correlation tensors, $C_{i\;j}^{\mathbf{u}\mathbf{u}}({\bf r},t)=\left\langle u_{i}u_{j}^{\prime}\right\rangle=\left\langle u_{i}({\bf x},t)u_{j}({\bf x}+{\bf r},t)\right\rangle$ and $C_{i\;j}^{\mathbf{h}\mathbf{h}}({\bf r},t)=\left\langle h_{i}h_{j}^{\prime}\right\rangle=\left\langle h_{i}({\bf x},t)h_{j}({\bf x}+{\bf r},t)\right\rangle$, which can be derived from the MHD equations (under the assumption of homogeneity, see Appendix B and C) according to
@@
$A_{[ki]j}^{\mathbf{u}\mathbf{h}{\mathbf{h}}}({\bf r},t)=$ $\left\langle(h_{i}u_{k}-h_{k}u_{i})h_{j}^{\prime}\right\rangle$
@@
Furthermore, the indices $r$ and $t$ denote the longitudinal and transverse projections, ${\bf u}_{r}=\frac{{\bf r}}{r}\left(\frac{{\bf r}}{r}\cdot{\bf u}\right)$ and ${\bf u}_{t}=-\left(\frac{{\bf r}}{r}\times\left(\frac{{\bf r}}{r}\times{\bf u}\right)\right)$ respectively (the same projections hold for the magnetic field ${\bf h}({\bf x},t)$, see Appendix F for further derivation). Following Chandrasekhar Chandrasekhar (1951) (see Appendix C for further derivations), we obtain evolution equations for the longitudinal velocity and magnetic field correlation functions
@@
$\frac{\partial}{\partial t}C_{r\;r}^{\mathbf{h}\mathbf{h}}(r,t)=$ $-\frac{4}{r}A_{[rt]t}^{\mathbf{u}\mathbf{h}\mathbf{h}}(r,t)+2\lambda\frac{1}{r^{4}}\frac{\partial}{\partial r}r^{4}\frac{\partial}{\partial r}C_{r\;r}^{\mathbf{h}\mathbf{h}}(r,t)\;.$ (9)
Eq. (8) is the generalization of the von Kármán-Howarth equation to MHD turbulence and the additional third-order correlation is due to the Lorentz force $C_{r\;r\;r}^{\mathbf{h}\mathbf{h}\mathbf{u}}(r,t)$ in the MHD equations. A similar equation also holds for the second-order longitudinal magnetic field correlation function (9) where the third-order correlation is due to the advection by the velocity field and the corresponding transformation of the electric field ${\bf E}^{\prime}={\bf E}+\mu{\bf u}\times{\bf H}$. Similar equations were derived recently in the context of Hall MHD Galtier (2008), where the Hall term modifies such a transformation.
@@
$\left.\left.+\frac{\partial}{\partial r}\left[\nu S_{r\;r}^{\mathbf{u}\mathbf{u}}(r,t)+\lambda S_{r\;r}^{\mathbf{h}\mathbf{h}}(r,t)\right]\right)\right]+\frac{4}{r}A_{[rt]t}^{\mathbf{u}\mathbf{h}\mathbf{h}}(r,t)\;.$ (17)
@@
In the inertial range, i.e., for $r$ smaller than integral- and larger than dissipation-length scales, we neglect the viscous terms in square brackets and obtain a relation between third-order statistics on the l.h.s. and a term that is proportional to $r$ on the r.h.s. whose magnitude is determined by the total averaged local energy dissipation rate $\langle\varepsilon_{tot}\rangle$ Friedrich et al. (2016). Nonetheless, in contrast to the hydrodynamic case (3) and to the original derivation by Politano and Pouquet Politano and Pouquet (1998a, b) where the energy transfer is purely local in scale $r$, we obtain a scaling law in the MHD inertial range (18) that exhibits a nonlocal term stemming from the magnetic induction equation and is a direct consequence of the anti-symmetric tensorial form (7). As shown in Appendix E, we can re-cast this nonlocal term as
@@
which thus represents the influence of magnetic field structures induced at larger scales $r^{\prime}$ on the local fluctuations at scale $r$ in the sense of the Alfvén effect Alfvén (1950); Biskamp (2003); Iroshnikov (1964); Kraichnan (1965). We will now consider further implications of this additional nonlocality for the conservation of energy at small scales and the phenomenon of solar wind heating by turbulent cascades. In more detail, for small scales $r$ in the vicinity of the dissipation range, nonlinear transfer terms on the l.h.s. in Eq. (18) should decay faster than the terms on the r.h.s. Monin and Yaglom (2007). From a Taylor expansion around $r=0$, we obtain $S_{r\;r\,r}^{\mathbf{u}\mathbf{u}\mathbf{u}}(r)\sim r^{3}$ as well as
@@
Figure 3: (a) Assessment of the four-fifths law in MHD turbulence from DNS ($1024^{3}$) of forced MHD turbulence simulation. The dashed line represents the r.h.s. of Eq. (18, which is determined from the total energy dissipation rate $\varepsilon_{tot}$ of the snapshot. The top curve corresponds the original P&P law $S_{r\;r\;r}^{\mathbf{u}\mathbf{u}\mathbf{u}}(r)-6C_{r\;r\;r}^{\mathbf{h}\mathbf{h}\mathbf{u}}(r)$ derived in Politano and Pouquet (1998a, a) whereas the additional nonlocal term in Eq. (18) is closer to the dashed line. The mixed tensors verify the dissipation range behavior (22). The lowest curve represents the third-order longitudinal velocity structure function $S_{r\;r\;r}^{\mathbf{u}\mathbf{u}\mathbf{u}}(r)$. (b) Same as in (a) but compensated by the r.h.s. in Eq. (18). The original P&P law (top curve) overestimates the total energy dissipation rate $\varepsilon_{tot}$ by a factor of 1.71 and peaks in front of the inertial range, whereas the inclusion of the nonlocal term in Eq. (18) leads to the correct prediction (dashed line).
@@
$\left.\frac{\partial A_{[rt]r}^{\mathbf{u}\mathbf{h}\mathbf{h}}(r)}{\partial r}\right|_{r=0}=-\frac{5}{4}\left.\frac{\partial C_{r\;r\;r}^{\mathbf{h}\mathbf{h}\mathbf{u}}(r)}{\partial r}\right|_{r=0}\;,$ (22)
@@ III Verification of the third-order law via direct numerical simulations of MHD turbulence
To verify Eq. (18), we performed DNS of homogeneous and isotropic MHD turbulence in a box of $1024^{3}$ grid points with characteristic turbulence parameters summarized in Tab. 1. The forcing scheme consists of a divergence-free random forcing Alvelius (1999) that is applied to the evolution equation of the velocity field (i.e., the evolution equation for the magnetic field is not actively forced). As shown in Fig. 2, after an initial phase, the forcing scheme leads to nearly constant kinetic and magnetic energy dissipation rates $\langle\varepsilon_{kin}\rangle$ and $\langle\varepsilon_{mag}\rangle$ over time. Fig. 3 depicts the third-order moments that enter Eq. (18). The lowest curve corresponds to the third-order velocity structure function $S_{r\;r\;r}^{\mathbf{u}\mathbf{u}\mathbf{u}}(r)$. Interestingly, this curve exhibits three distinct scaling regions: In the dissipation range, we observe the predicted scaling $\sim r^{3}$ followed by a region that scales in agreement with the Iroshnikov-Kraichnan phenomenology $\sim r^{3/4}$ whereas a linear increase seems to dominate at larger $r$. Furthermore, in the dissipation range, we can directly verify the relation between the symmetric and anti-symmetric correlations (22). In the case under study, the original prediction Politano and Pouquet (1998a, b) without the additional nonlocal source term overestimates the total energy dissipation rate $\langle\varepsilon_{tot}\rangle$ (dashed line) and does not vanish for small $r$, therefore, violating energy conservation. This becomes even more apparent in the compensated plot in Fig. 3 (b), which suggests that including the source term results in the correct scaling in the inertial range in between the largest Taylor scale $\lambda_{kin}$ and the smallest integral length scale $L_{mag}$. In contrast, the scaling obtained using the P&P law Politano and Pouquet (1998a, b) (green curve) would plateau before this inertial range and overestimate the $\langle\varepsilon_{tot}\rangle$ by a factor of 1.71. We must emphasize that this discrepancy between the original P&P law and (18) might have been barely detectable in previous numerical studies of MHD turbulence Mininni and Pouquet (2009); Sorriso-Valvo et al. (2002); Yousef et al. (2007) due for instance to low magnetic and fluid Reynolds numbers and signatures of the large-scale anisotropic forcing scheme (e.g., by Taylor-Green vortices). Indeed, MHD and kinetic simulations reproduce plasma dynamics and features peculiar to the set of governing parameters defined for specific runs, which may differ significantly from actual solar wind parameters. Due to the insufficient computational power of today’s supercomputers, constraints in reproducing the solar wind numerically are technical but also inherent since solar wind dynamics develop in a vast parameter space, hardly reproducible even with large ensembles of simulations.